Calculate the kinetic energy in units of electron volts for an electron emitted from a surface with a work function of 1.5 eV that has been illuminated with632.8-nm (red) helium neon laser, such as those used in a supermarket scanner. Perform the same calculation for a 543.5-nm (green) helium-neon laser.
Answer #1

Kinetic energy of electron emitted from red helium neon laser will be:

KE = E - W

E = energy of wavelength = h*c/\lambda

E = 6.626*10^-34*3*10^8/(632.8*10^-9) = 3.14*10^-19 J

E = (3.14*10^-19)/(1.6*10^-19) eV = 1.96 eV

W = Work-function = 1.5 eV

So,

KE_red = 1.96 eV - 1.5 eV = 0.46 eV

For green light:

E = energy of wavelength = h*c/\lambda

E = 6.626*10^-34*3*10^8/(543.5*10^-9) = 3.657*10^-19 J

E = (3.657*10^-19)/(1.6*10^-19) eV = 2.29 eV

So,

KE_green = 2.29 - 1.5 = 0.79 eV

Please Upvote.